Please explain and solve this for me:
(-14) - (-128) =
and
(-12) - (-7) =
and
What is the commutative property of addition?
What is the associative property of multiplication?
What is an additive inverse?

THANKS !

Answer by CubeyThePenguin(3113)

(-14) - (-128) = 114
(-12) - (-7) = -5

commutative property of addition: a + b = b + a

associative property of muliplication: a(bc) = (ab)(c)

additive inverse: The additive inverse of a is -a. A number added to its additive inverse results in 0.